OpenStudy (anonymous):

a tugboat goes 120 miles upstream is 15hours. the return trip downstream takes 5 hours.find the speed of the tugboat without a current and the speed of the current.

OpenStudy (anonymous):

Not 100% sure, ask @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Hint: Let x = speed of the boat (without the current) and c = speed of the current Upstream equation: d = rt 120 = (x-c)*15 -------------------- Downstream equation: d = rt 120 = (x+c)*5

jimthompson5910 (jim_thompson5910):

You have 2 equations with 2 unknowns, which you can solve for both x and c.

OpenStudy (anonymous):

the speed of the tugboat is blank mph, and the speed of the current is blank mph.. what is both speed

jimthompson5910 (jim_thompson5910):

120 = (x-c)*15 120/15 = x-c 8 = x-c x = 8 + c ------------------------- 120 = (x+c)*5 120 = (8+c+c)*5 120 = (8+2c)*5 120 = 40 + 10c Now solve for c